Two Coast Gaurd cutters are 225 miles apart traveling toward each other, one from the east and the other from the west. The eastbound cutter travels 5 miles per hour faster than the westbound cutter. If the two cutters pass each other after 3 hours, find the average speed of each cutter.
Answer by mananth(16949)

Two Coast Gaurd cutters are 225 miles apart traveling toward each other, one from the east and the other from the west. The eastbound cutter travels 5 miles per hour faster than the westbound cutter. If the two cutters pass each other after 3 hours, find the average speed of each cutter.
..
west bound speed be x
eastbound speed = x+5
...
they are travelling towards each other.
so effective speed = 2x+5 mph.
..
distance = 225 miles
time = 3 hours
distance/speed = time
225/2x+1= 3
225 = 3(2x+1)
225 = 6x+3
225-3 = 6x
222=6x
/6
37 = x
westbound speed = 37 mph
eastbound is 5 mph more
...
m.ananth@hotmail.ca